### v1

@@ Abstract
Mitral regurgitation is one of the most prevalent cardiac disorders. Four-dimensional (4D) ultrasound has emerged as the primary imaging modality for assessing dynamic valvular morphology. However, 4D mitral valve (MV) analysis remains challenging due to limited phase annotations, severe motion artifacts, and poor imaging quality. Yet, the absence of inter-phase dependency in existing methods hinders 4D MV analysis. To bridge this gap, we propose a Motion-Topology guided consistency network (MTCNet) for accurate 4D MV ultrasound segmentation in semi-supervised learning (SSL). MTCNet requires only sparse end-diastolic and end-systolic annotations. First, we design a cross-phase motion-guided consistency learning strategy, utilizing a bi-directional attention memory bank to propagate spatio-temporal features. This enables MTCNet to achieve excellent performance both per- and inter-phase. Second, we devise a novel topology-guided correlation regularization that explores physical prior knowledge to maintain anatomically plausible. Therefore, MTCNet can effectively leverage structural correspondence between labeled and unlabeled phases. Extensive evaluations on the first largest 4D MV dataset, with 1408 phases from 160 patients, show that MTCNet performs superior cross-phase consistency compared to other advanced methods (Dice: 87.30%, HD: 1.75mm). Both the code and the dataset are available at https://github.com/crs524/MTCNet.
@@ 1 Introduction
Despite the promising progress in deep learning [26, 13, 9, 8], research on 4D MV segmentation remains limited, with most studies focusing on single-volume segmentation. Some studies [3, 1] employed U-Net [19] for 3D MV segmentation, while others [5, 17] utilized the nnU-Net [10] framework. Although these methods achieve promising accuracy in static 3D volumes, they struggle with MV motion due to a lack of information extraction from unannotated phases. Semi-supervised learning (SSL) [20, 27, 28] has emerged as a promising paradigm to enhance model performance with limited labeled data by utilizing the abundant information in unlabeled data. Ivantsits et al. [11] simplify MV as an idea tubular sheet to reconstruct 4D MV surface, which limits its use in simulating the common organic mitral regurgitation disease with valve abnormalities. Munaf‘o et al. [18] proposed an SSL method for 4D MV segmentation, where ED and ES phases of all patients are treated as labeled, while intermediate phases are considered unlabeled ones. However, they treat each phase independently, neglecting patient-level phase consistency, leading to suboptimal segmentation.
@@ 2.2 Motion-guided Consistency Learning
To effectively enhance the inter-phase motion coherence, we propose an MCL strategy through the BMB block. As shown in Fig. 2 (a), we propose forward memory bank $M_{f}$ and backward memory bank $M_{b}$ to store the temporal information for all phases of a patient. $M_{f}$ captures systolic-diastolic deformation, while $M_{b}$ encodes reverse motion, jointly modeling bidirectional patterns to reduce phase misalignment in cardiac cycles.
Fig. 3 illustrates the detailed design of the BMB block. Here, we take a memory bank to illustrate the mechanism, as both follow the same process. Given a memory bank $M$ containing $T$ memory phases, the memory encoder would generate memory key $k^{M}\in\mathbb{R}^{C^{k}\times TDHW}$ and memory value $v^{M}\in\mathbb{R}^{C^{v}\times TDHW}$. Similarly, the query encoder produces a query key $k^{Q}\in\mathbb{R}^{C^{k}\times DHW}$ and a query value $k^{Q}\in\mathbb{R}^{C^{k}\times DHW}$, where $D$, $H$, and $W$ are the multi-scale feature dimensions. Current phase feature is the output feature of encoders both in the student and teacher models. Therefore, multi-scale memory aggregation minimizes information loss, especially for subtle boundary changes in low signal-to-noise ratio ultrasound phases. Then, we compute the normalized affinity matrix $\mathbf{W}$ to to effectively weigh the dependencies across phases:
@@
With the normalized affinity matrix $\mathbf{W}$, the aggregated readout feature $v^{Q}\in\mathbb{R}^{C^{v}\times DHW}$ for the query phase is computed as a weighted sum of the memory features using a top-k operation. To be specific, the weighted sum of the top-k memory features is calculated as: $v_{r}^{Q}=v_{r}^{M}\mathbf{W}_{r}$, where $r$ represents two forward and backward memory banks. Finally, the $v_{1}^{Q}$ and $v_{2}^{Q}$ are concatenated to obtain the bi-directional attention feature. This feature is stored in the memory banks $M_{f}$ and $M_{b}$ and passed to the decoder to produce the segmentation mask.
@@ 2.3 Topology-guided Consistency Regulation
The dual-term design overcomes key limitations of static shape constraints. The relative term adapts to patient-specific anatomy via the annotated $S(P_{1})$, while the absolute term prevents error propagation through physical consistency. Follow a similar strategy to surface area consistency, the volume consistency loss can be formulated based on the voxel-wise summation of $B_{t}$:
@@ 3 Experiments and Results
Datasets and Evaluation Metrics. We collected 4D TEE MV data from 160 patients with 1408 phases from cooperating hospitals. The in-house dataset comprised 147 cases with functional mitral regurgitation, and 13 normal cases. The dataset was randomly split into training (112 cases), validation (16 cases), and testing (32 cases) subsets. The number of phases varied per case due to heart rate differences, with an average of nine phases per cycle. All imaging data were acquired using the Philips X5-1 transthoracic volume probe. Two sonographers with more than 10 years of experience manually annotated the MV volumes, consisting anterior leaflets (AL) and posterior leaflets (PL). We also validated on the mid-diastolic (MD) phase and its adjacent transitional phase (MD-1) between MD and ED. For evaluation, Dice coefficient, 95th-percentile Hausdorff Distance (HD) and Conformity (Conf) [4], were adopted for quantitative comparison.
Comparison Study. We evaluated MTCNet against multiple SSL methods [20, 27, 12, 14, 23, 24, 7], the registration-based approach VoxelMorph (VM) [2], and the two-shot video object segmentation method (T-VOS) [25]. VM and T-VOS both require one first phase as reference. As depicted in Table 1, MTCNet achieves state-of-the-art (SOTA) performance, surpassing compared methods in Dice ($87.30\%$) and Conf ($66.71\%$) in all phases segmentation. Specifically, our method surpasses the best baseline ICT (86.23%) by 1.07% (p$<$0.05). Notably, although MD is the most challenging phase due to the leaflets are maximally separated and often near the left ventricular wall, MTCNet surpasses the sub-optimal T-VOS method by 2.4% ($82.69\%$ vs. $80.29\%$) in Dice, highlighting MTCNet ’s ability in handling complex deformation. Additionally, compared to reference-phase-driven methods VM and T-VOS, MTCNet shows more potential results in capturing continuous motion more effectively. Moreover, MTCNet achieves robust performance for both labeled phases (ES and ED) and unlabeled phases (MD and MD-1). This highlights MTCNet ’s superior ability in per-phase and inter-phase performance.
@@
Table 2: Ablation study of MTCNet. M and T represent MCL and TCR, respectively. Four key phases denote the four key phases (ES, MD, MD-1, ED). * denotes statistically significant differences (t-test) between best model and other configurations.
@@
Qualitative Results. We further perform a qualitative comparison. Fig. 4 clearly shows that MTCNet produces smoother and more complete segmentations. Particularly in the MD phase where severe motion and deformation often cause holes (bule arrows). Artifacts in this region are most prominent due to the valve’s motion, indicating MTCNet ’s superior ability to handle complex motion and structure changes. To further validate the clinical applicability of our approach, we compared the 3D printing results (See Fig. 5). While other methods produce incomplete shapes and missing details, our model achieves greater realism and completeness, showcasing its potential for patient-specific planning.
@@ 4 Conclusion
In this study, we present MTCNet, a SSL framework that leverages motion and topology-guided consistency learning for 4D MV segmentation. MTCNet performs better than the SOTA with sparse annotation, especially in motion and structure coherence. We contribute this gain in performance in two key contributions. First, the proposed MCL strategy can learn motion information across phases through BMB mechanism, improving segmentation performance in per-phase and inter-phase. Additionally, a robust TCR strategy with surface and volume variance as prior knowledge boosts segmentation accuracy while preserving topological coherence. MTCNet can also facilitate MV dynamics analysis and support personalized hemodynamic assessment and treatment planning.

### v2

@@ Abstract
Mitral regurgitation is one of the most prevalent cardiac disorders. Four-dimensional (4D) ultrasound has emerged as the primary imaging modality for assessing dynamic valvular morphology. However, 4D mitral valve (MV) analysis remains challenging due to limited phase annotations, severe motion artifacts, and poor imaging quality. Yet, the absence of inter-phase dependency in existing methods hinders 4D MV analysis. To bridge this gap, we propose a Motion-Topology guided consistency network (MTCNet) for accurate 4D MV ultrasound segmentation in semi-supervised learning (SSL). MTCNet requires only sparse end-diastolic and end-systolic annotations. First, we design a cross-phase motion-guided consistency learning strategy, utilizing a bi-directional attention memory bank to propagate spatio-temporal features. This enables MTCNet to achieve excellent performance both per- and inter-phase. Second, we devise a novel topology-guided correlation regularization that explores physical prior knowledge to maintain anatomical plausibility. Therefore, MTCNet can effectively leverage structural correspondence between labeled and unlabeled phases. Extensive evaluations on the first largest 4D MV dataset, with 1408 phases from 160 patients, show that MTCNet performs superior cross-phase consistency compared to other advanced methods (Dice: 87.30%, HD: 1.75mm). Both the code and the dataset are available at https://github.com/crs524/MTCNet.
@@ 1 Introduction
Despite the promising progress in deep learning [26, 13, 9, 8], research on 4D MV segmentation remains limited, with most studies focusing on single-volume segmentation. Some studies [3, 1] employed U-Net [19] for 3D MV segmentation, while others [5, 17] utilized the nnU-Net [10] framework. Although these methods achieve promising accuracy in static 3D volumes, they struggle with MV motion due to a lack of information extraction from unannotated phases. Semi-supervised learning (SSL) [20, 27, 28] has emerged as a promising paradigm to enhance model performance with limited labeled data by utilizing the abundant information in unlabeled data. Ivantsits et al. [11] simplify MV as an ideal tubular sheet to reconstruct a 4D MV surface, which limits its use in simulating the common organic mitral regurgitation disease with valve abnormalities. Munaf‘o et al. [18] proposed an SSL method for 4D MV segmentation, where ED and ES phases of all patients are treated as labeled, while intermediate phases are considered unlabeled ones. However, they treat each phase independently, neglecting patient-level phase consistency, leading to suboptimal segmentation.
@@ 2.2 Motion-guided Consistency Learning
To effectively enhance the inter-phase motion coherence, we propose an MCL strategy through the BMB block. As shown in Fig. 2 (a), we propose forward memory bank $M_{f}$ and backward memory bank $M_{b}$ to store the temporal information for all phases of a patient. $M_{f}$ captures systolic-diastolic deformation, while $M_{b}$ encodes reverse motion, jointly modeling bi-directional patterns to reduce phase misalignment in cardiac cycles.
Fig. 3 illustrates the detailed design of the BMB block. Here, we take a memory bank to illustrate the mechanism, as both follow the same process. Given a memory bank $M$ containing $T$ memory phases, the memory encoder would generate memory key $k^{M}\in\mathbb{R}^{C^{k}\times TDHW}$ and memory value $v^{M}\in\mathbb{R}^{C^{v}\times TDHW}$. Similarly, the query encoder produces a query key $k^{Q}\in\mathbb{R}^{C^{k}\times DHW}$ and a query value $v^{Q}\in\mathbb{R}^{C^{v}\times DHW}$, where $D$, $H$, and $W$ are the multi-scale feature dimensions. The current phase feature is the output feature of encoders, both in the student and teacher models. Therefore, multi-scale memory aggregation minimizes information loss, especially for subtle boundary changes in low signal-to-noise ratio ultrasound phases. Then, we compute the normalized affinity matrix $\mathbf{W}$ to effectively weigh the dependencies across phases:
@@
With the normalized affinity matrix $\mathbf{W}$, the aggregated readout feature $v^{Q}\in\mathbb{R}^{C^{v}\times DHW}$ for the query phase is computed as a weighted sum of the memory features using a top-k operation. To be specific, the weighted sum of the top-k memory features is calculated as: $v_{r}^{Q}=v_{r}^{M}\mathbf{W}_{r}$, where $r$ represents two forward and backward memory banks. Finally, the $v_{f}^{Q}$ and $v_{b}^{Q}$ are concatenated to obtain the bi-directional attention feature. This feature is stored in the memory banks $M_{f}$ and $M_{b}$ and passed to the decoder to produce the segmentation mask.
@@ 2.3 Topology-guided Consistency Regulation
The dual-term design overcomes key limitations of static shape constraints. The relative term adapts to patient-specific anatomy via the annotated $S(P_{1})$, while the absolute term prevents error propagation through physical consistency. Following a similar strategy to surface area consistency, the volume consistency loss can be formulated based on the voxel-wise summation of $B_{t}$:
@@ 3 Experiments and Results
Datasets and Evaluation Metrics. We collected 4D TEE MV data from 160 patients with 1408 phases from cooperating hospitals. The in-house dataset comprised 147 cases with functional mitral regurgitation, and 13 normal cases. The dataset was randomly split into training (112 cases), validation (16 cases), and testing (32 cases) subsets. The number of phases varied per case due to heart rate differences, with an average of nine phases per cycle. All imaging data were acquired using the Philips X5-1 transthoracic volume probe. Two sonographers with more than 10 years of experience manually annotated the MV volumes, consisting of anterior leaflets (AL) and posterior leaflets (PL). We also validated the mid-diastolic (MD) phase and its adjacent transitional phase (MD-1) between MD and ED. For evaluation, Dice coefficient, 95th-percentile Hausdorff Distance (HD), and Conformity (Conf) [4], were adopted for quantitative comparison.
Comparison Study. We evaluated MTCNet against multiple SSL methods [20, 27, 12, 14, 23, 24, 7], the registration-based approach VoxelMorph (VM) [2], and the two-shot video object segmentation method (T-VOS) [25]. VM and T-VOS both require the first phase as reference. As depicted in Table 1, MTCNet achieves state-of-the-art (SOTA) performance, surpassing compared methods in Dice ($87.30\%$) and Conf ($66.71\%$) in all phases of segmentation. Specifically, our method surpasses the best baseline ICT (86.23%) by 1.07% (p$<$0.05). Notably, although MD is the most challenging phase due to the leaflets are maximally separated and often near the left ventricular wall, MTCNet surpasses the sub-optimal T-VOS method by 2.4% ($82.69\%$ vs. $80.29\%$) in Dice, highlighting MTCNet ’s ability in handling complex deformation. Additionally, compared to reference-phase-driven methods VM and T-VOS, MTCNet shows more potential results in capturing continuous motion more effectively. Moreover, MTCNet achieves robust performance for both labeled phases (ES and ED) and unlabeled phases (MD and MD-1). This highlights MTCNet ’s superior ability in per-phase and inter-phase performance.
@@
Table 2: Ablation study of MTCNet. M and T represent MCL and TCR, respectively. Four key phases denote the four key phases (ES, MD, MD-1, ED). * denotes statistically significant differences (t-test) between the best model and other configurations.
@@
Qualitative Results. We further perform a qualitative comparison. Fig. 4 clearly shows that MTCNet produces smoother and more complete segmentations. Particularly in the MD phase, where severe motion and deformation often cause holes (bule arrows). Artifacts in this region are most prominent due to the valve’s motion, indicating MTCNet ’s superior ability to handle complex motion and structure changes. To further validate the clinical applicability of our approach, we compared the 3D printing results (See Fig. 5). While other methods produce incomplete shapes and missing details, our model achieves greater realism and completeness, showcasing its potential for patient-specific planning.
@@ 4 Conclusion
In this study, we present MTCNet, an SSL framework that leverages motion and topology-guided consistency learning for 4D MV segmentation. MTCNet performs better than the SOTA with sparse annotation, especially in motion and structure coherence. We attribute this gain in performance to two key contributions. First, the proposed MCL strategy can learn motion information across phases through BMB mechanism, improving segmentation performance in per-phase and inter-phase. Additionally, a robust TCR strategy with surface and volume variance as prior knowledge boosts segmentation accuracy while preserving topological coherence. MTCNet can also facilitate MV dynamics analysis and support personalized hemodynamic assessment and treatment planning.